OpenStudy (anonymous):

Slope7; through (-7,5) i need to write the equation of the line

OpenStudy (chaise):

y=mx+c y=7x+c y=5 x=-7 5=7*-7+c 5=-49+c c=54 y=7x+54 I hope this is right.

OpenStudy (anonymous):

you just need to find the y-intercept now. plug in the given point and slope into y =mx + b where y= y coord, x= x coor, and m= slope. y =mx +b 5 = (7)(-7)+ b 5 = -49 + b 54 = b

OpenStudy (anonymous):

so the equation is y = 7x + 54 ^oh, @chaise is correct :)

OpenStudy (anonymous):

The equation of a Line with Slope given and passing through a point is given by: \[y - y_1 = m(x - x_1)\] Where m is the Slope, (x, y) are the Arbitrary Points and (x1, y1) are the given points through which the line is passing..